OpenStudy (anonymous):

Solve the equation by finding square roots. 36x^2=256

OpenStudy (anonymous):

36X^2=256 or 9x^2=64 9x^2-64=0 or (3x)^2-(8)^2=0 or(3x+8)(3x-8)=0 or 3x+8=0 or 3x-8=0 or x=-8/3 or x=8/3